OpenStudy (anonymous):

3-27x^3=0 is this solvable?

OpenStudy (lgbasallote):

anything is solvable for a mathematician. they believe in imaginary things

OpenStudy (anonymous):

Yep, it is: \[ 3-27x^3=0\implies\\ 3=27x^3 \]Taking the third root of both sides: \[ \sqrt[3]{3}=\sqrt[3]{27x^3}=3x\implies\\ x=\frac{\sqrt[3]{3}}{3} \]This is for real \(x\).

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Sure thing!